# Learning to chain-of-thought with Jensen’s evidence lower bound

### v1

@@ Abstract
We propose a way to optimize chain-of-thought with reinforcement learning, but without external reward function. Our algorithm relies on viewing chain-of-thought as latent variable as part of a probabilistic inference problem. Contrary to the full evidence lower bound, we propose to apply a much simpler Jensen’s lower bound, which derives tractable objectives with simple algorithmic components (e.g., without the need for parametric approximate posterior), making it more conducive to modern large-scale training. The lower bound approach naturally interpolates other methods such as supervised fine-tuning and online reinforcement learning, whose practical trade-offs we will illustrate. Finally, we show that on mathematical reasoning problems, optimizing with Jensen’s lower bound is as effective as policy gradient with external reward. Taken together, our results showcase as a proof of concept to this new algorithmic paradigm’s potential to more generic applications.
@@
1 Introduction
Chain-of-thought is playing an increasingly important role for large language models. It has been observed that chain-of-thoughts can elicit significant performance improvements from language models, which scales positively with model sizes (Wei et al., 2022). Chain-of-thought can be instantiated in different ways, in the case of reasoning, it usually refers to the step-by-step solution leading up to a final conclusion (Ling et al., 2017; Cobbe et al., 2021; Lightman et al., 2023). Learning to reason with chain-of-thought has proved highly effective, and can endow models with capabilities well beyond pre-training and regular fine-tuning (Jaech et al., 2024; Guo et al., 2025; Lambert et al., 2024; Team et al., 2025).
There are many algorithmic approaches to optimizing chain-of-thoughts for language models, such as prompting, supervised fine-tuning (SFT) and reinforcement learning (RL) (Ling et al., 2017; Cobbe et al., 2021; Wei et al., 2022; Jaech et al., 2024). A rather intriguing perspective is to view chain-of-thought optimization as a latent variable modeling problem. In short, given an initial problem, chain-of-thoughts are interpreted as intermediary latent variables such that the desirable conclusion becomes more likely (Hoffman et al., 2024; Hu et al., 2024). Such an interpretation naturally fits the intuition we have for chain-of-thoughts, that they should serve as intermediary outputs that improve the final prediction. This formulation also brings about the algorithmic developments from the latent variable modeling literature.
In this work, we take a pragmatic approach to adapting methodologies from the latent variable modeling literature. We propose an algorithm based on a single-sample or multiple-samples version of the Jensen’s evidence lower bound, named after the fact that our objective is obtained by simply applying a Jensen’s inequality to the evidence (Jensen, 1906). Optimizing with such a simple lower bound forgoes the complication of variational posterior, which is inherent to the full evidence lower bound (Blei et al., 2017; Kingma and Welling, 2013b; Rezende et al., 2014). This makes our algorithm more suitable for contemporary large-scale training for language model applications, where training auxiliary models is as expensive as the main model (Brown et al., 2020; Achiam et al., 2023).
The final algorithm takes a rather simple form - in a nutshell, it consists of an hybrid online RL-like and SFT-like loss. When viewed as an alternative to online RL, the lower bound method does not require any external reward, unlike regular RLHF algorithms (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022). In more details, our technical contributions are as follows:
  • We derive the Jensen’s evidence lower bound in Section 3, the stochastic optimization algorithm, as well as its multi-sample extension (Burda et al., 2015) which tightens the bound in Section 4.
  • We draw insightful connections between the full ELBO, RL with specialized reward and the proposed lower bound algorithm in Section 5. We also show how regularized RLHF as a whole can be coherently interpreted as a latent variable modeling problem. See Figure 1 for a graphical illustration.
  • Besides theoretical equivalence, we highlight the practical trade-offs of applying RL vs. lower bound optimization. The key trade-off factors include the choice of rewards, which we detail in Section 6.
  • Finally in Section 8, we show that optimizing with Jensen’s evidence lower bound and its multi-sample variants is competitive compared to online RL, even without explicit access to a more sophisticated reward function.
2 Reinforcement learning for language model
with a KL regularization that encourages $\pi_{\theta}$ to stay close to the reference policy. The reward $r(x,y)$ captures the human preference of response $y$ in response to prompt $x$ and can take various forms: for example, it can be extracted from human annotations (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022), generated atuomatic feedback such as code execution (Gehring et al., 2024). We focus on a specialized setting where the reward is derived from access to certain ground truth of the problem.
2.1 RL from ground truth feedback
We focus on applications where the prompt $x$ typically specifies a question and there is an example of an desirable ground truth $a^{\ast}$. Such a formulation includes formally verifiable applications such as mathematical reasoning (Hendrycks et al., 2021; Uesato et al., 2022; Lightman et al., 2023) where $x$ is a question and $a^{\ast}$ is the ground truth answer. Usually, the reward $r(x,y)$ is calculated by matching $a^{\ast}$ against the answer $a$, e.g., using string match.
2.2 Chain-of-thought
For aforementioned applications where the model is required to reason about the question $x$ and generate an answer $a$, getting the model to generate chain-of-thoughts - a sequence of reasoning steps $c$ leading up to the final conclusion (Ling et al., 2017; Wei et al., 2022). Henceforth, we can decompose the generation $y=(c,a)$ into a chain-of-thought $c$ and an answer $a$. The generative process for the response $y\sim\pi_{\theta}(\cdot|x)$ is made more concrete as
$c\sim\pi_{\theta}(\cdot|x),a\sim\pi_{\theta}(\cdot|x,c).$ (2)
3 A Jensen’s lower bound for chain-of-thought as latent variable modeling
@@
$\log\pi_{\theta}(a^{\ast}|x)-\mathcal{L}_{\theta}(x,a^{\ast})=\mathbb{KL}\left(\pi_{\theta}(\cdot|x),p^{\pi_{\theta}}(\cdot|x,a^{\ast})\right),$
where $p^{\pi_{\theta}}(c|x,a^{\ast})\coloneqq\frac{\pi_{\theta}(a^{\ast}|x,c)\pi_{\theta}(c|x)}{\sum_{c^{\prime}}\pi_{\theta}(a^{\ast}|x,c^{\prime})\pi_{\theta}(c^{\prime}|x)}$. The posterior defines a distribution over chain-of-thought, and effectively denotes how likely is the chain-of-thought $c$ given that the ground truth answer is $a=a^{\ast}$ and the prompt is $x$. For experienced readers, this lower bound is closely related to the evidence lower bound (Kingma and Welling, 2013a; Blei et al., 2017), which we will elaborate more below.
@@ 3.1 Stochastic gradient estimate
In practice, we can add a control variate to the REINFORCE gradient estimate to reduce variance. One option is to learn a prompt-answer dependent function (Schulman et al., 2017); another sample-based alternative is to generate $n$ i.i.d. chain-of-thoughts in parallel $c_{i}\sim\pi_{\theta}(\cdot|x)$, and construct leave-one-out control variates $v_{i}=\frac{1}{n-1}\sum_{j\neq i}\log\pi_{\theta}(a^{\ast}|x,c_{j})$ (Mnih and Rezende, 2016; Kool et al., 2019). The overall gradient estimate is the average over $n$ samples:
$\frac{1}{n}\sum_{i=1}^{n}\Big{[}\left(\log\pi_{\theta}(a^{\ast}|x,c_{i})-v_{i}\right)\nabla_{\theta}\log\pi_{\theta}(c_{i}|x)$
$+\nabla_{\theta}\log\pi_{\theta}(a^{\ast}|x,c_{i})\Big{]}.$
Note the control variates $v_{i}$s do not introduce any bias to the gradient estimate since they are statistically independent from $\nabla_{\theta}\log\pi_{\theta}(c_{i}|x)$.
Connections to supervised fine-tuning.
In the very special case where there is no chain-of-thought, the gradient estimate reduces to just the SFT part $\nabla_{\theta}\log\pi_{\theta}(a^{\ast}|x)$ which is effectively the supervised fine-tuning loss from prompt $x$ to answer $a^{\ast}$. Here, the key difference is that the loss $\pi_{\theta}(a^{*}|x,c_{i})$ further conditions on the chain-of-thoughts $c_{i}$s whose distribution changes over time and introduce more non-stationarity to the optimization process.
4 Tightening the objective via multi-sample Jensen’s lower bound
If the Jensen’s lower bound is loose, it will induce a sizable discrepancy from the true objective of interest. We need strategies to tighten the lower bound for policy optimization.
A similarly simple yet a tighter lower bound alternative, is an extension to the multi-sample case (Burda et al., 2015). Indeed, consider the $n$-sample Jensen’s lower bound
@@ 4 Tightening the objective via multi-sample Jensen’s lower bound
Importantly, the $\log$ function is outside of the $n$-sample average to tighten the bound. It is straightforward to verify that $\mathcal{L}_{\theta}^{(1)}(x,a^{\ast})$ recovers the Jensen’s lower bound as defined before in Eqn (4). As shown in Burda et al. (2015), the lower bound becomes tighter as $n$ increases $\mathcal{L}_{\theta}^{(n)}(x,a^{\ast})\leq\mathcal{L}_{\theta}^{(n+1)}(x,a^{\ast})$ for any $n\geq 0$. As $n\rightarrow\infty$, the bound approaches the marginal likelihood $\mathcal{L}_{\theta}^{(n)}(x,a^{\ast})\rightarrow\log\pi_{\theta}(a^{\ast}|x)$, the ultimate objective of interest, under certain regularity conditions on $\pi_{\theta}$.
To maximize the multi-sample lower bound $\mathcal{L}_{\theta}^{(n)}(x,a^{*})$ with gradient ascent, we can construct the REINFORCE stochastic gradient estimate as follows,
@@
Empirically, the first term $g_{1}^{(n)}$ tends to have high variance as $n$ increases (Rainforth et al., 2018), since the objective $\log\frac{1}{n}\sum_{j=1}^{n}\pi_{\theta}(a^{\ast}|x,c_{j})$ correlates updates to all $n$ samples. Akin to before, we can introduce the leave-one-out control variate without incurring any bias for variance reduction (Mnih and Rezende, 2016; Kool et al., 2019)
$\sum_{i=1}^{n}\left(\log\left(\frac{1}{n}\sum_{j=1}^{n}\pi_{\theta}(a^{\ast}|x,c_{j})\right)-\tilde{v}_{i}\right)\cdot\nabla_{\theta}\log\pi_{\theta}(c_{i}|x)$
where $\tilde{v}_{i}=\log\frac{1}{n-1}\sum_{j\neq i}\pi_{\theta}(a^{\ast}|x,c_{j})$.
@@
Note that the second term $g_{2}^{(n)}$, though can be estimated via random samples, is unlike a regular SFT loss since it is the log average of multiple probabilities, instead of the average of log probabilities. As $n\rightarrow\infty$, since $\log\frac{1}{n}\sum_{i=1}^{n}\pi_{\theta}(a^{*}|x,c_{i})\rightarrow\log\pi_{\theta}(a^{*}|x)$, we see that at least conceptually $g_{2}^{(n)}$ can be understood as directly maximizing the marginal likelihood - the average over probabilities effectively marginalize the chain-of-thought conditional distribution.
@@
As we will show in Section 8, multi-sample lower bound generally improves the single sample Jensen’s lower bound. This means that tightened lower bound improve training objectives both in theory and in practice.

### v2

# Beyond Verifiable Rewards: Scaling Reinforcement Learning for Language Models to Unverifiable Data

@@ Abstract
We propose to scale RL to unverifiable data with a novel algorithm JEPO (Jensen’s Evidence lower bound Policy Optimization). While most prior efforts on scaling RL for LLMs focus on verifiable data where ground truth answers are typically short-form and can be matched easily; we investigate the case where such assumptions are less valid (e.g., when answers are long-form such as mathematical proofs). To scale RL training to unverifiable data with contemporary training constraints, we propose JEPO. JEPO applies Jensen’s evidence lower bound, a pragmatic simplification of the evidence lower bound which views chain-of-thought as a latent variable in the generative process. We show that on verifiable data (math), JEPO is as effective as RL with verifiable rewards; on semi-verifiable data (numina), JEPO improves on soft-match based evaluations compared to RL with verifiable rewards which can only leverage a subset of the data source; finally, on unverifiable data (numina-proof), JEPO outperforms SFT and a few ablation baselines on likelihood evaluations.
@@
1 Introduction
Reinforcement learning from verifiable rewards (RLVR) has proved effective at endowing language models with capabilities beyond canonical pre-training and supervised fine-tuning (Jaech et al., 2024; Shao et al., 2024; Lambert et al., 2024; Guo et al., 2025; Team et al., 2025; Su et al., 2025). At its core, reinforcement learning (RL) allows for the optimization of chain-of-thought at scale, which elicits significant performance improvements especially for reasoning intensive tasks (Ling et al., 2017; Wei et al., 2022). In the case of mathematical reasoning, it encourages step-by-step solutions that lead up to a final answer (Cobbe et al., 2021; Lightman et al., 2023), where correctness can be verified to produce a reward signal for RL training.
However, a main limitation of current RLVR is the data source: verifiable rewards are mostly derived from datasets where ground truth answers are short-form and can be checked in relatively easy ways (Guo et al., 2025; Team et al., 2025; Su et al., 2025). For example, most answers to popular benchmarks are integers and short expressions (Hendrycks et al., 2021; AoPS, 1983). This practical limitation makes it hard to scale RL to more general datasets where answer correctness is hard to check. For instance, for long-form mathematical data where the answer is the whole proof, its inherent correctness is hard to assess without expert human evaluations (Petrov et al., 2025).
Figure 1: A canonical RL algorithm updates both its chain-of-thought policy $\pi_{\theta}(c|x)$ and the final conclusion $\pi_{\theta}(a|x,c)$ with advantage function computed from reward $r_{i}$ and an optional baseline $v_{i}$. JEPO has similar counterparts: updating the chain-of-thought policy using likelihood scores as the effective reward, and updating the answer policy using a supervised loss. Unlike RL baselines, JEPO does not require access to a reward $r_{i}$ but only access to a ground truth answer $a^{*}$. Due to the implementation-level similarity between JEPO and RL, it is straightforward to incorporate JEPO into existing stacks of large-scale RL training. We use the same baseline notation for the RL and JEPO loss, though they differ in practice. In general $v_{i}$ can be a leave-one-out control variate that is computed from other $n-1$ samples in the batch.
The boundary between verifiable and unverfiable data, though often blurry in practice, can be made more actionable: we define data as unverifiable, if its ground truth answer cannot be verified with a reasonably simple automatic procedure. Naturally, it is of interest to scale RL to such data sources, for a few notable reasons: (1) some data have inherently long answers which cannot be cast into short-form answers in a straightforward way; (2) data sources with long-form answers exist in abundance, and it is sub-optimal not to leverage such data for training. In this work, we seek to tackle the problem of scaling RL to unverifiable data.
We propose JEPO (Jensen’s Evidence lower bound Policy Optimization), a novel RL algorithm that can equally post-train on verifiable or unverifiable data. The design of the algorithm is inspired by a latent variable view of chain-of-thought (Hoffman et al., 2024; Hu et al., 2024). As a major algorithmic innovation, contrast to prior work, we make use of Jensen’s evidence lower bound, a novel pragmatic simplification of the full evidence lower bound (Blei and Jordan, 2006; Blei et al., 2017) named after Jensen’s inequality (Jensen, 1906). Optimizing such a simplified objective forgoes the need of training expensive auxiliary models, making JEPO more suitable for contemporary large-scale training (Brown et al., 2020; Achiam et al., 2023).
The final algorithm consists of an hybrid RL and supervised learning loss. As a major advantage over online RL baselines, JEPO does not require any external verifiable reward, lifting the requirement that ground truth be easily verifiable. JEPO also shares much implementation-level similarity with online RL algorithms, making it easy to integrate into an existing large-scale workflow. See Figure 1 for a visual depiction of the similarity and difference between JEPO and RL baselines. In more details, our technical contributions are as follows:
  • (Algorithm) Followed by a brief background on latent variable modeling, we derive the Jensen’s evidence lower bound in Section 3. In Section 4, we show how its multi-sample extension (Burda et al., 2015) tightens the theoretical bound and alludes to better performance in practice. For all objectives, we derive stochastic optimization algorithms that can be practically implemented.
  • (Theoretical connections) We draw insightful connections between the full ELBO, RL and JEPO in Section 5. We discuss a few different connections of interest to readers from different backgrounds, such as the practical trade-offs of RL vs. JEPO. See Figure 2 for an illustration of the graphical models connecting JEPO and probabilistic inference.
  • (Implementation) In Section 6, we highlight practical implementation details that make JEPO work the best, highlighting the fact that the resulting algorithm takes a similar form to common RL algorithms for LLM. This means that JEPO is easy to integrate into an existing workflow. See Figure 1 for a summarized comparison.
  • (Experiments) Finally in Section 8, Section 9 and Section 10, we show that for verifiable data, JEPO is competitive compared to online RL with verifiable reward. For semi-verifiable and unverifiable data, JEPO has performance advantage over online RL, SFT or other ablation baselines. As a by-product, we showcase the utility of generating chain-of-thought for long-form proofs, an observation that is interesting in its own right.
2 Reinforcement learning for language models
with a KL regularization that encourages $\pi_{\theta}$ to stay close to the reference policy. The reward $r(x,y)$ captures the human preference of response $y$ in response to prompt $x$ and can take various forms: for example, it can be extracted from human annotations (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022), computed using automatic feedback such as code execution (Gehring et al., 2024; Wei et al., 2025). We focus on a specialized setting where the reward is derived from access to a certain ground truth of the problem.
2.1 RL from ground truth feedback
We focus on applications where the prompt $x$ typically specifies a question and there is an example of a desirable ground truth $a^{\ast}$. Such a formulation is applicable to mathematical reasoning (Hendrycks et al., 2021; Uesato et al., 2022; Lightman et al., 2023) where $x$ is a question and $a^{\ast}$ is the ground truth answer. When the correctness of the model generated answer $a$ can be easily verified against the ground truth $a^{*}$, a verifiable reward $r$ is available by matching $a^{\ast}$ against the answer $a$. As another example, when $a^{*}$ is a long-form proof, such a reward is not immediately available and such cases are considered less verifiable.
In broader context, RLVR also includes code applications where the reward is computed via unit tests (Gehring et al., 2024; Wei et al., 2025). We do not consider such use cases.
2.2 Chain-of-thought
For aforementioned applications where the model is required to reason about the question $x$ and generate an answer $a$, we get the model to generate chain-of-thoughts - a sequence of reasoning steps $c$ leading up to the final conclusion (Ling et al., 2017; Wei et al., 2022). Henceforth, we can decompose the generation $y=(c,a)$ into a chain-of-thought $c$ and an answer $a$. The generative process for the response $y\sim\pi_{\theta}(\cdot|x)$ is made more concrete as
$c\sim\pi_{\theta}(\cdot|x),a\sim\pi_{\theta}(\cdot|x,c).$ (2)
@@ 2.2 Chain-of-thought
3 Jensen’s lower bound for chain-of-thought as latent variable modeling
@@
$\log\pi_{\theta}(a^{\ast}|x)-\mathcal{L}_{\theta}(x,a^{\ast})=\mathbb{KL}\left(\pi_{\theta}(\cdot|x),p^{\pi_{\theta}}(\cdot|x,a^{\ast})\right)\geq 0,$
where $p^{\pi_{\theta}}(c|x,a^{\ast})\coloneqq\frac{\pi_{\theta}(a^{\ast}|x,c)\pi_{\theta}(c|x)}{\sum_{c^{\prime}}\pi_{\theta}(a^{\ast}|x,c^{\prime})\pi_{\theta}(c^{\prime}|x)}$ is the posterior, which defines a distribution over chain-of-thought given the prior $\pi_{\theta}(c|x)$ and the likelihood $\pi_{\theta}(a^{*}|x,c)$. For readers familiar with the probabilistic inference literature. The lower bound $\mathcal{L}_{\theta}(x,a^{*})$ is closely related to the evidence lower bound (Kingma and Welling, 2013; Blei et al., 2017), which we will elaborate more in Section 5.
@@ 3.1 Stochastic gradient estimate
In practice, we can add a control variate to the REINFORCE gradient estimate to reduce variance. One option is to learn a prompt-answer dependent function (Schulman et al., 2017); another sample-based alternative is to generate $n$ i.i.d. chain-of-thoughts in parallel $c_{i}\sim\pi_{\theta}(\cdot|x)$, and construct leave-one-out control variates $v_{i}=\frac{1}{n-1}\sum_{j\neq i}\log\pi_{\theta}(a^{\ast}|x,c_{j})$ (Mnih and Rezende, 2016; Kool et al., 2019; Tang et al., 2025). The overall gradient estimate is the average over $n$ samples:
$\frac{1}{n}\sum_{i=1}^{n}\Big{[}\left(\log\pi_{\theta}(a^{\ast}|x,c_{i})-v_{i}\right)\nabla_{\theta}\log\pi_{\theta}(c_{i}|x)\Big{]}$
$+\frac{1}{n}\sum_{i=1}^{n}\Big{[}\nabla_{\theta}\log\pi_{\theta}(a^{\ast}|x,c_{i})\Big{]}.$ (6)
Note the control variates $v_{i}$s do not introduce any bias to the gradient estimate since they are statistically independent from $\nabla_{\theta}\log\pi_{\theta}(c_{i}|x)$ and $\log\pi_{\theta}(a^{*}|x,c_{i})$.
Connections to supervised fine-tuning
In the very special case where there is no chain-of-thought, the gradient estimate reduces to just the SFT part $\nabla_{\theta}\log\pi_{\theta}(a^{\ast}|x)$ which is effectively the supervised fine-tuning loss from prompt $x$ to answer $a^{\ast}$. Here, the key difference is that the loss $\pi_{\theta}(a^{*}|x,c_{i})$ further conditions on the chain-of-thoughts $c_{i}$’s whose distribution changes over time and introduces more diversity to the optimization process.
4 Improving the objective via multi-sample Jensen’s lower bound
A loose lower bound induces a sizable discrepancy from the true objective of interest. A similarly simple yet tighter lower bound can be obtained with multiple samples (Burda et al., 2015). Indeed, consider the $n$-sample lower bound
@@ 4 Improving the objective via multi-sample Jensen’s lower bound
Note that the $\log$ function is outside of the $n$-sample average to tighten the bound. It is straightforward to verify that $\mathcal{L}_{\theta}^{(1)}(x,a^{\ast})$ recovers the Jensen’s lower bound as defined before in Eqn (4). As shown in Burda et al. (2015), the lower bound becomes tighter as $n$ increases $\mathcal{L}_{\theta}^{(n)}(x,a^{\ast})\leq\mathcal{L}_{\theta}^{(n+1)}(x,a^{\ast})$ for any $n\geq 0$. As $n\rightarrow\infty$, the bound approaches the marginal likelihood $\mathcal{L}_{\theta}^{(n)}(x,a^{\ast})\rightarrow\log\pi_{\theta}(a^{\ast}|x)$, which is the ultimate objective of interest, under certain regularity conditions on $\pi_{\theta}$.
To maximize the multi-sample lower bound $\mathcal{L}_{\theta}^{(n)}(x,a^{*})$ with gradient ascent, we can construct a multi-sample stochastic gradient estimate as follows,
@@
Empirically, the first term $g_{1}^{(n)}$ tends to have high variance as $n$ increases (Rainforth et al., 2018), since the objective $\log\frac{1}{n}\sum_{j=1}^{n}\pi_{\theta}(a^{\ast}|x,c_{j})$ correlates updates to all $n$ samples. As a result, a key difference from the single-sample case is that the update is no longer an average over $n$ samples (Tang et al., 2025). Akin to before, we can introduce the leave-one-out control variate without incurring any bias for variance reduction (Mnih and Rezende, 2016; Kool et al., 2019) with $\tilde{v}_{i}=\log\frac{1}{n-1}\sum_{j\neq i}\pi_{\theta}(a^{\ast}|x,c_{j})$,
$\sum_{i=1}^{n}\left(\log\left(\frac{1}{n}\sum_{j=1}^{n}\pi_{\theta}(a^{\ast}|x,c_{j})\right)-\tilde{v}_{i}\right)\cdot\nabla_{\theta}\log\pi_{\theta}(c_{i}|x).$
@@
Note that the second term $g_{2}^{(n)}$, though can be estimated via random samples, is unlike a regular SFT loss. The key difference is that it is the log average of multiple probabilities, instead of the average of log probabilities as in the regular SFT loss. As $n\rightarrow\infty$, since $\log\frac{1}{n}\sum_{i=1}^{n}\pi_{\theta}(a^{*}|x,c_{i})\rightarrow\log\pi_{\theta}(a^{*}|x)$, we see that conceptually $g_{2}^{(n)}$ can be understood as directly maximizing the marginal likelihood. In other words, the objective averages over multiple probabilities, which essentially marginalizes the chain-of-thought conditional distribution.
@@
As we will show in Section 8, multi-sample lower bound generally improves the single-sample lower bound. This means that tightened lower bounds improve training objectives both in theory and in practice.
5 Connections to algorithmic alternatives
The lower bound objectives bear close connections to a number of algorithmic alternatives, which we discuss below. See Algorithm 1 for the pseudocode of the full algorithm, which we henceforth call JEPO.
5.1 Evidence lower bound
The evidence lower bounds (ELBO) (Blei and Jordan, 2006; Kingma and Welling, 2013; Burda et al., 2015) controls for the tightness of the lower bound with an inference distribution $q_{\phi}(c|x,a^{\ast})$ which defines a distribution over chain-of-thoughts. ELBO is usually written as follows
$\mathcal{L}_{\theta,\phi}(x,a^{\ast})=\mathbb{E}_{c}\left[\log\pi_{\theta}(a^{\ast}|x,c)-\log\frac{q_{\phi}(c|x,a^{*})}{\pi_{\theta}(c|x)}\right],$ (9)
where the expectation is under $c\sim q_{\phi}(\cdot|x,a^{\ast})$. ELBO lower bounds the marginal log likelihood $\mathcal{L}_{\theta,\phi}(x,a^{\ast})\leq\log\pi_{\theta}(a^{\ast}|x)$ and it is tight if and only if the inference distribution equals the posterior distribution $q_{\phi}(c|x,a^{\ast})=p^{\pi_{\theta}}(c|x,a^{\ast})$. Since ELBO is a function of both the policy parameter $\theta$ and inference distribution parameter $\phi$, given a chain-of-thought sample $c\sim q_{\phi}(\cdot|x,a^{\ast})$, we can optimize both with stochastic gradient estimates:
$\ \ -\nabla_{\phi}\log q_{\phi}(c|x,a^{*}).$
Juxtaposing the form of the gradient here and the gradient to the Jensen’s lower bound defined in Eqn (5), we observe that the inference distribution gradient $g_{\phi}$ bears resemblance to the REINFORCE gradient; while the policy distribution gradient $g_{\theta}$ bears resemblance to the SFT gradient. In fact, we can show that under the special parameterization $q_{\phi}(c|x,a^{\ast})\coloneqq\pi_{\theta}(c|x)$, the two gradients are exactly equivalent. More formally, we have the following.
Lemma 1.
(Jensen’s lower bound as a special case of ELBO) When $q_{\phi}(c|x,a^{\ast})\coloneqq\pi_{\theta}(c|x)$, ELBO is equivalent to the Jensen’s lower bound $\mathcal{L}_{\theta,\phi}(x,a^{\ast})=\mathcal{L}_{\theta}(x,a^{\ast})$ stochastic gradient estimates.
Proof.
When $q_{\phi}=\pi_{\theta}$, we have
With a parametric approximate posterior $q_{\phi}$, ELBO is more expressive than the Jensen’s lower bound and allows for a tighter approximation to the marginal log likelihood. However, this also introduces additional complexity of having to learn the approximate posterior distribution. In our applications of interest, training a posterior model of a large size can be a major computational overhead. In practice, for example, Hoffman et al. (2024) approximates the posterior via a few steps of MCMC and avoids learning such a distribution. We take a different approach with a similar motivation: by tightening the lower bound with multiple samples, we also avoid the need for a parametric approximate posterior.
5.2 Reinforcement learning
We show that there is a close connection between the lower bound formulation and the expected return maximization objective in RL (Sutton and Barto, 1998) for a single terminal reward. Concretely, we will see how the lower bound objectives are closely related to a conditional expectation trick that produces a RL policy gradient estimate with lower variance. First, we show that (up to a log transform) RL optimizes for the same target as the lower bound objectives, given the indicator reward.
Lemma 2.
Proof.